Show that the total interest paid on the amortization of a loan with leveled monthly payments K for n years at an interest rate of i% per annum is equal


1
Expert's answer
2021-08-01T16:16:40-0400

Monthly payments=k per month

Number of years=n years

Interest rate="\frac{i}{100}" Per annum

Let C be the total amount of loan after n years

"=K.12.n"

Let P be initial payment (loan amount)

"P=\frac{K[1-(1+\frac{i}{12×100})^{-12×n}]}{\frac{i}{12×100}}"


Therefore,

Total interest paid="C-P"

"=K.12.n-\frac{K[1-(1+\frac{i}{1200})^{-12n}]}{\frac{i}{1200}}"

"=\frac{K.12n.\frac{i}{100}-K+K(1+\frac{i}{1200})^{-12n}]}{\frac{i}{1200}}"

"=\frac{K(\frac{i}{100}-1)+K(1+\frac{i}{1200})^{-12n}}{\frac{i}{1200}}"

"=\frac{1200K[(\frac{i}{100}-1)+(1+\frac{i}{1200})^{-12n}]}{i}"